• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Glesener

    Log entry time 01:16:49 on July23,2005

    Entry number 149085

    keyword=CHL/ HPH trouble

    The flow from CHL dropped, raising the temperature and pressure of the target.  
    The pressure again rose above 230 psi.  The high power heater went to zero.  We 
    turned the beam off and conditions returned to normal.  The drop in CHL was 
    temporary-it has also returned to normal.  We turned off the beam compensation 
    because of this and the earlier trouble with the heater.
    


    FIGURE 1

    FIGURE 2